Rusty Nails Bar & Grill reported free cash flow of $779,000.00 last year. Free cash flows are projected to grow by 2.50% annually. For the same period, the firm reported marketable securities of $9,100.00 and $700,000.00 in outstanding long-term debt. The firm’s reported shareholder equity totaled $300,000.00. The firm currently has 100,000.00 common shares outstanding and the cost of capital is 13.00%.

1) What is the value of operations for Rusty Nails Bar & Grill?

2) What is the total corporate (firm) value for Rusty Nails Bar & Grill?

3) What is the intrinsic value of common equity for Rusty Nails Bar & Grill?

4) What is the intrinsic value of common equity per share for Rusty Nails Bar & Grill?
Answer #1

1)Value of operations =Free cash Flow*(1+growth)/Cost of Capital-Growth)=779000*(1+2.5%)/(13%-2.5%) =7604523.81
2)Total Corporate Value = Value of Equity+Marketable securities+Debt =300000+9100+700000 =1009100
3)Intrinsic value = Value of operations -Marketable securities-Debt =7604523.81-9100-700000=6895423.81
4)Value per share =Intrinsic value/Number of shares =6895423.81/100000 =68.95